# Scalable Graph Attention‐Based Instance Selection via Mini‐Batch Sampling and Hierarchical Hashing

### v1

@@ Abstract
Instance selection (IS) is important in machine learning for reducing dataset size while keeping key characteristics. Current IS methods often struggle with capturing complex relationships in high-dimensional spaces and scale with large datasets. This paper introduces a graph attention-based instance selection (GAIS) method that uses attention mechanisms to identify informative instances through their structural relationships in graph representations. We present two approaches for scalable graph construction: a distance-based mini-batch sampling technique that reduces computation through strategic batch processing, and a hierarchical hashing approach that allows for efficient similarity computation through random projections. The mini-batch approach keeps class distributions through stratified sampling, while the hierarchical hashing method captures relationships at multiple granularities through single-level, multi-level, and multi-view variants. Experiments across 39 datasets show that GAIS achieves reduction rates above 96% while maintaining or improving model performance relative to state-of-the-art IS methods. The findings shows that the distance-based mini-batch approach offers an optimal balance of efficiency and effectiveness for large-scale datasets, while multi-view variants provide superior performance for complex, high-dimensional data, demonstrating that attention-based importance scoring can effectively identify instances crucial for maintaining decision boundaries without requiring exhaustive pairwise comparisons.
@@ 1 Introduction
  • Development of GAIS (Graph Attention-based Instance Selection), a novel framework that leverages attention mechanisms to identify informative instances in tabular data
  • Introduction of mini-batch sampling approach that reduces graph construction complexity from $O(n^{2})$ to $O(n^{2}/K)$ while preserving class distributions
  • Design of multi-level and multi-view LSH variants that capture relationships at different granularities through random projections
  • Extensive empirical evaluation demonstrating significant improvements over state-of-the-art IS methods across diverse datasets, with particular effectiveness in complex, high-dimensional problems
@@ 2.3 Scaling Challenges in Graph-based Methods
Figure 3: High-level Overview of the Methodology.
Beyond graph construction, computational challenges persist in the subsequent model training phase. Training GNNs on dense graphs requires substantial computational resources for message passing between nodes. Additionally, dense graphs can result the over-smoothing problem in GNNs (Rusch et al., 2023), where repeated message passing causes node representations to converge to similar values. This over-smoothing effect leads to a loss of discriminative power after multiple layers, potentially undermining the GNN’s ability to identify truly informative instances for selection.
@@ 3.1.1 Distance-based Mini-Batching Method
Figure 4: High-level overview of the mini-batch distance metric-based GAIS.
@@
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}$ is the corresponding class label, we partition $\mathcal{D}$ into $K$ mini-batches $\{B_{1},\ldots,B_{K}\}$ using stratified sampling without replacement to preserve class distribution. Computing pairwise distances within each $B_{k}$ reduces the overall complexity from $O(n^{2})$ to approximately $O(n^{2}/K)$ when mini-batches are balanced. The choice of $K$ balances efficiency and model performance; larger $K$ offers finer processing but increases overhead.
@@
where $w_{\max}$ is the maximum batch size and $f_{\text{original}}$ is the original sampling fraction. This ensures $|B_{k}|\leq w_{\max}$ by reducing the sampling fraction whenever $f_{\text{original}}\cdot n$ would exceed $w_{\max}$. It also preserves the class distribution $P(y=c\,\mid\,x\in B_{k})\approx P(y=c\,\mid\,x\in\mathcal{D})$ for all classes $c$.
@@
Computationally, this approach reduces the naive $O(n^{2})$ distance-computation cost to $O\left(\sum_{k=1}^{K}|B_{k}|^{2}\right).$ Under balanced partitioning where $|B_{k}|\approx{n}/{K}$, the cost approximates $O(n^{2}/K)$. This methodology leverages stochastic approximation principles and stratified random sampling, creating mini-batches that reflect the dataset’s underlying distribution and leading to more reliable, generalizable results while keeping computations feasible.
@@ 3.1.2 Locality-Sensitive Hashing-based Method
Figure 5: High-level overview of the Multi-Level LSH approach. Lower levels capture coarse similarities with fewer hash tables and projections, while higher levels employ exponential growth in hashing parameters to refine local relationships.
@@ 3.2.3 Diversity-Promoting Attention
The final attention weights incorporate this diversity term:
@@ 4.4 Hyperparameter Optimization
Figure 6: High-level overview of the hyperparameter tuning procedure.
@@ 4.5 Evaluation Metrics
Table 2: Evaluation Metrics
@@ 5 Results and Discussions
Table 3: Summary of performance metrics for GAIS methods across different dataset sizes.
@@
SL-HLSH 0.813 ± 0.169 0.705 ± 0.249 0.680 ± 0.280 0.682 ± 0.260 0.786 ± 0.170 0.659 ± 0.254 0.966 ± 0.023
ML-HLSH 0.823 ± 0.174 0.726 ± 0.257 0.675 ± 0.288 0.686 ± 0.270 0.792 ± 0.173 0.659 ± 0.260 0.961 ± 0.029
CML-HLSH 0.823 ± 0.170 0.720 ± 0.255 0.674 ± 0.288 0.684 ± 0.268 0.793 ± 0.170 0.657 ± 0.259 0.961 ± 0.025
MVML-HLSH 0.824 ± 0.173 0.730 ± 0.256 0.682 ± 0.294 0.689 ± 0.273 0.794 ± 0.171 0.662 ± 0.263 0.963 ± 0.026
CMVML-HLSH 0.825 ± 0.171 0.745 ± 0.240 0.696 ± 0.285 0.704 ± 0.264 0.796 ± 0.171 0.678 ± 0.255 0.963 ± 0.025
@@
SL-HLSH 0.895 ± 0.051 0.816 ± 0.160 0.781 ± 0.181 0.783 ± 0.168 0.865 ± 0.063 0.755 ± 0.160 0.966 ± 0.031
ML-HLSH 0.894 ± 0.054 0.810 ± 0.148 0.789 ± 0.184 0.789 ± 0.166 0.866 ± 0.061 0.762 ± 0.157 0.968 ± 0.024
CML-HLSH 0.896 ± 0.053 0.812 ± 0.146 0.787 ± 0.184 0.788 ± 0.167 0.869 ± 0.062 0.762 ± 0.158 0.969 ± 0.025
MVML-HLSH 0.895 ± 0.054 0.810 ± 0.152 0.784 ± 0.183 0.790 ± 0.166 0.868 ± 0.062 0.764 ± 0.155 0.970 ± 0.021
CMVML-HLSH 0.893 ± 0.057 0.809 ± 0.155 0.776 ± 0.192 0.781 ± 0.171 0.864 ± 0.062 0.754 ± 0.159 0.968 ± 0.020
@@
SL-HLSH 0.823 ± 0.111 0.685 ± 0.233 0.516 ± 0.296 0.547 ± 0.291 0.799 ± 0.123 0.531 ± 0.288 0.969 ± 0.032
ML-HLSH 0.824 ± 0.111 0.706 ± 0.221 0.516 ± 0.299 0.548 ± 0.295 0.798 ± 0.124 0.532 ± 0.291 0.966 ± 0.034
CML-HLSH 0.823 ± 0.112 0.692 ± 0.234 0.523 ± 0.283 0.558 ± 0.276 0.800 ± 0.123 0.544 ± 0.274 0.971 ± 0.029
MVML-HLSH 0.840 ± 0.109 0.696 ± 0.208 0.477 ± 0.294 0.519 ± 0.280 0.819 ± 0.123 0.506 ± 0.277 0.973 ± 0.031
CMVML-HLSH 0.823 ± 0.114 0.715 ± 0.219 0.528 ± 0.292 0.563 ± 0.287 0.800 ± 0.122 0.547 ± 0.283 0.970 ± 0.029
Table 4: Summary of computational time metrics (in seconds) for GAIS methods across different dataset sizes.
@@ 5.1 Comparing GAIS Variants
In medium-sized datasets, the performance gap between the original and GAIS narrows considerably. The original achieves an accuracy of 0.945, while MVML-LSH maintains robust performance with an accuracy of 0.895 and effectiveness of 0.868. The DM method demonstrates stable performance (E = 0.872) in medium datasets, outperforming several LSH variants, which suggests that its adaptive similarity threshold mechanism scales well with increasing data size. Interestingly, the multi-view and cross-attention variants show marginal differences in effectiveness (ranging from 0.864 to 0.869), suggesting that the additional complexity of cross-attention mechanisms might not provide substantial benefits. The reduction rates remain consistently high across all proposed methods (96-97%).
For large datasets, while MVML-HLSH achieves the highest effectiveness (E = 0.819) and accuracy (AC = 0.840), its cross-attention variant CMVML-HLSH demonstrates superior performance across other crucial metrics, including precision (PR = 0.715), recall (RE = 0.528) and F1-score (F1 = 0.563). This suggests that the cross-attention mechanism enhances the model’s ability to identify more informative instances, leading to better balanced performance. The DM approach shows competitive effectiveness (E = 0.795) while maintaining a high reduction rate (R = 0.98), demonstrating its viability as a computationally efficient alternative. The random sampling baselines consistently underperform compared to the proposed methods across all metrics, validating the importance of structured selection approaches.
@@
Table 4 presents a summary of computational time metrics for GAIS methods across different dataset sizes, comparing graph construction time (GCT), GAT (or GNN) training time (GTT), selection time (IST), and ML training time (MLT). For small datasets, all proposed methods demonstrate efficient computational performance, with GCT remaining under 2 seconds on average. The DM method shows particularly efficient graph construction (GCT = 0.05s), making it an attractive option for time-sensitive applications with small datasets. While the multi-level variants (ML-LSH, CML-LSH) show slightly higher GCT ($\approx$ 1.9s), they maintain competitive GTT around 4-5 seconds. Notably, the IST is negligible across all methods for small datasets.
@@ 5.2 Comparing Graph Timings
Figure 7(a) compares graph construction times between GAIS variants and KNN-based approaches across medium and large datasets. The DM method consistently demonstrates higher efficiency across all dataset sizes, maintaining relatively stable construction times even as dataset size increases. Among the LSH variants, SL-LSH shows the best computational efficiency, followed by ML-LSH, while MVML-LSH exhibits higher construction times due to its more complex multi-view architecture. When compared to KNN-based approaches, all GAIS variants show better scalability, particularly for larger datasets. This is most evident in datasets like SUSY, where exact KNN and FAISS exhibit exponential growth in construction time, while GAIS methods maintain more controlled growth. Notably, while approximate KNN methods like Annoy and HNSW show better scaling than exact KNN, they still require significantly more time than the GAIS variants, especially for datasets exceeding 100K instances. These results demonstrate that the proposed GAIS methods, particularly the DM approach, offer a more scalable alternative to traditional KNN-based graph construction methods.
@@ 5.3 Component Analysis
Figure 8 illustrates the relationship between mini-batch configuration and test effectiveness in the DM method for large datasets. The results demonstrate that both the number of mini-batches and instances per batch have a positive impact on effectiveness, though this improvement shows diminishing returns. As the number of mini-batches increases from 1 to 9, the effectiveness consistently improves across all instance sizes, with the most substantial gains observed in the transition from 1 to 5 mini-batches. Similarly, larger batch sizes generally yield better effectiveness, though the improvement plateaus around 7000 instances, particularly with higher numbers of mini-batches. This plateau effect is most noticeable when using 7-9 mini-batches, suggesting that beyond these configurations, additional computational resources might not yield proportional performance benefits. These findings provide practical guidance for implementing the DM method, indicating that moderate configurations (5-7 mini-batches with 5000-7000 instances per batch) might offer an optimal balance between computational efficiency and effectiveness.
Figure 8: Impact of mini-batch configuration on test effectiveness in the DM method for large datasets.

### v2

@@ Abstract
Instance selection (IS) addresses the critical challenge of reducing dataset size while keeping informative characteristics, becoming increasingly important as datasets grow to millions of instances. Current IS methods often struggle with capturing complex relationships in high-dimensional spaces and scale with large datasets. This paper introduces a graph attention-based instance selection (GAIS) method that uses attention mechanisms to identify informative instances through their structural relationships in graph representations. We present two approaches for scalable graph construction: a distance-based mini-batch sampling technique that achieves dataset-size-independent complexity through strategic batch processing, and a hierarchical hashing approach that enables efficient similarity computation through random projections. The mini-batch approach keeps class distributions through stratified sampling, while the hierarchical hashing method captures relationships at multiple granularities through single-level, multi-level, and multi-view variants. Experiments across 39 datasets show that GAIS achieves reduction rates above 96% while maintaining or improving model performance relative to state-of-the-art IS methods. The findings show that the distance-based mini-batch approach offers an optimal efficiency for large-scale datasets, while multi-view variants excel on complex, high-dimensional data, demonstrating that attention-based importance scoring can effectively identify instances important for maintaining decision boundaries while avoiding computationally prohibitive pairwise comparisons.
@@ 1 Introduction
  • Development of GAIS (Graph Attention-based Instance Selection), a novel framework that uses attention mechanisms to identify informative instances in tabular data.
  • Introduction of mini-batch sampling approach that reduces graph construction complexity from $O(n^{2})$ to $O(K\times w_{\max}^{2})$ while preserving class distributions through strategic sampling with computational bounds.
  • Design of multi-level and multi-view LSH variants that capture relationships at different granularities through random projections.
  • Extensive empirical evaluation demonstrating significant improvements over state-of-the-art IS methods across diverse datasets, with particular effectiveness in complex, high-dimensional problems.
@@ 2.3 Scaling Challenges in Graph-based Methods
Figure 3: High-level overview of the Graph Attention-based Instance Selection (GAIS) methodology. The framework operates in three stages: (1) Graph Construction - creates scalable graph representations using either distance-based mini-batch sampling that selects instances without replacement to form subgraphs, or Locality-Sensitive Hashing with single-level, multi-level, and multi-view variants that use all instances; (2) Graph Attention Network - applies basic GAT for mini-batching/single-level LSH and enhanced multi-level/multi-view GAT with cross-attention via Nyström approximation to learn node importance through attention weights; (3) Instance Selection - computes importance scores from attention weights and applies either top-percentage global selection or class-aware selection (balanced or proportional) to retain the most informative instances.
Beyond graph construction, computational challenges persist in the subsequent model training phase. Training GNNs on dense graphs requires substantial computational resources for message passing between nodes. Also, dense graphs can result the over-smoothing problem in GNNs (Rusch et al., 2023), where repeated message passing causes node representations to converge to similar values. This over-smoothing effect leads to a loss of discriminative power after multiple layers, potentially undermining the GNN’s ability to identify truly informative instances for selection.
@@ 3.1.1 Distance-based Mini-Batching Method
Figure 4: High-level overview of the mini-batch distance metric-based (DM) GAIS. The method samples $K$ mini-batches from the dataset $\mathcal{D}$ without replacement while preserving class distributions and ensuring each batch contains at most $w_{\max}$ instances. For each mini-batch, we compute pairwise distances to construct dense subgraphs, which are then sparsified using similarity thresholds to retain only the strongest connections. The resulting subgraphs are stacked and processed by a graph attention network (GAT). Lastly, attention weights from the trained GAT are aggregated across all $K$ mini-batches to compute node importance scores for selection.
@@
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}$ is the corresponding class label, we sample $K$ mini-batches $\{B_{1},\ldots,B_{K}\}$ from $\mathcal{D}$ using stratified sampling without replacement to preserve class distribution. Computing pairwise distances within each $B_{k}$ reduces the overall complexity from $O(n^{2})$ to $O(K\times w_{\max}^{2})$, where $w_{\max}$ is the maximum instances per mini-batch. The choice of $K$ balances efficiency and model performance; larger $K$ offers finer processing but increases overhead, while $w_{\max}$ provides computational bounds.
@@
where $f_{\text{original}}$ is the original sampling fraction. This ensures $|B_{k}|\leq w_{\max}$ by reducing the sampling fraction whenever $f_{\text{original}}\cdot n$ would exceed $w_{\max}$. It also preserves the class distribution $P(y=c\,\mid\,x\in B_{k})\approx P(y=c\,\mid\,x\in\mathcal{D})$ for all classes $c$. This constraint-based sampling approach differs from traditional partitioning by implementing strategic sampling with computational bounds. Rather than exhaustively dividing the dataset into $K$ equal parts, we sample $K$ mini-batches where each contains at most $w_{\max}$ instances. This ensures scalable complexity regardless of dataset size.
@@
Computationally, this approach reduces the naive $O(n^{2})$ distance-computation cost to $O\left(\sum_{k=1}^{K}|B_{k}|^{2}\right)=O(K\times w_{\max}^{2})$, since $|B_{k}|\leq w_{\max}$ for all mini-batches. This constraint ensures computational complexity remains independent of dataset size, making the approach truly scalable for large datasets. This methodology leverages stochastic approximation principles and stratified random sampling, creating mini-batches that reflect the dataset’s underlying distribution and leading to more reliable, generalizable results while keeping computations feasible. The $w_{\max}$ constraint is empirically justified through plateau analysis (see Component Analysis), where performance stabilizes at mini-batch sizes well within typical computational limits.
@@ 3.1.2 Locality-Sensitive Hashing-based Method
Figure 5: High-level overview of the Multi-Level LSH (ML-LSH) approach. ML-LSH processes instances through multiple hierarchical levels (Level $1$ to $M$), where each level follows a four-step process: (1) Random projections using hash functions, (2) Hash code generation (e.g., $h(x)=\text{sign}(P^{T}x)$), (3) Bucket assignment grouping similar instances, and (4) Edge connection within buckets. Lower levels (Level $1$) capture coarse relationships using fewer hash tables and projections, while higher levels (Level $M$) employ exponentially increasing hashing parameters to refine local similarities. The final graph combines edge sets from all levels, with each level treated independently to capture different relationship granularities.
@@ 3.2.3 Diversity-Promoting Attention
where $\epsilon$ is a small positive constant (typically $\epsilon=10^{-8}$) added for numerical stability to prevent division by zero when the mean attention $\bar{\alpha}_{i}^{k}$ approaches zero. The final attention weights incorporate this diversity term:
@@ 4.4 Hyperparameter Optimization
Figure 6: Hyperparameter tuning procedure using Bayesian optimization with nested trials. The main loop (left) conducts 50 trials optimizing graph construction and GAT parameters, while each trial includes a nested $\rho$-tuning subprocess (right) that performs 20 trials to optimize instance selection retention rates and strategies.
@@ 4.5 Evaluation Metrics
Table 2: Metrics used in GAIS evaluation. Classification metrics use standard definitions where TP=True Positives, TN=True Negatives, FP=False Positives, FN=False Negatives. For multi-class problems, precision, recall, and F1 are macro-averaged.
@@ 5 Results and Discussions
Table 3: Performance comparison of GAIS variants across dataset sizes. Methods include distance-based mini-batching (DM), locality-sensitive hashing variants (SL/ML/MVML-LSH with cross-attention C), and random sampling baselines. Values show mean $\pm$ standard deviation across multiple runs. Bold indicates best performance per dataset category.
@@
SL-LSH 0.813 ± 0.169 0.705 ± 0.249 0.680 ± 0.280 0.682 ± 0.260 0.786 ± 0.170 0.659 ± 0.254 0.966 ± 0.023
ML-LSH 0.823 ± 0.174 0.726 ± 0.257 0.675 ± 0.288 0.686 ± 0.270 0.792 ± 0.173 0.659 ± 0.260 0.961 ± 0.029
CML-LSH 0.823 ± 0.170 0.720 ± 0.255 0.674 ± 0.288 0.684 ± 0.268 0.793 ± 0.170 0.657 ± 0.259 0.961 ± 0.025
MVML-LSH 0.824 ± 0.173 0.730 ± 0.256 0.682 ± 0.294 0.689 ± 0.273 0.794 ± 0.171 0.662 ± 0.263 0.963 ± 0.026
CMVML-LSH 0.825 ± 0.171 0.745 ± 0.240 0.696 ± 0.285 0.704 ± 0.264 0.796 ± 0.171 0.678 ± 0.255 0.963 ± 0.025
@@
SL-LSH 0.895 ± 0.051 0.816 ± 0.160 0.781 ± 0.181 0.783 ± 0.168 0.865 ± 0.063 0.755 ± 0.160 0.966 ± 0.031
ML-LSH 0.894 ± 0.054 0.810 ± 0.148 0.789 ± 0.184 0.789 ± 0.166 0.866 ± 0.061 0.762 ± 0.157 0.968 ± 0.024
CML-LSH 0.896 ± 0.053 0.812 ± 0.146 0.787 ± 0.184 0.788 ± 0.167 0.869 ± 0.062 0.762 ± 0.158 0.969 ± 0.025
MVML-LSH 0.895 ± 0.054 0.810 ± 0.152 0.784 ± 0.183 0.790 ± 0.166 0.868 ± 0.062 0.764 ± 0.155 0.970 ± 0.021
CMVML-LSH 0.893 ± 0.057 0.809 ± 0.155 0.776 ± 0.192 0.781 ± 0.171 0.864 ± 0.062 0.754 ± 0.159 0.968 ± 0.020
@@
SL-LSH 0.823 ± 0.111 0.685 ± 0.233 0.516 ± 0.296 0.547 ± 0.291 0.799 ± 0.123 0.531 ± 0.288 0.969 ± 0.032
ML-LSH 0.824 ± 0.111 0.706 ± 0.221 0.516 ± 0.299 0.548 ± 0.295 0.798 ± 0.124 0.532 ± 0.291 0.966 ± 0.034
CML-LSH 0.823 ± 0.112 0.692 ± 0.234 0.523 ± 0.283 0.558 ± 0.276 0.800 ± 0.123 0.544 ± 0.274 0.971 ± 0.029
MVML-LSH 0.840 ± 0.109 0.696 ± 0.208 0.477 ± 0.294 0.519 ± 0.280 0.819 ± 0.123 0.506 ± 0.277 0.973 ± 0.031
CMVML-LSH 0.823 ± 0.114 0.715 ± 0.219 0.528 ± 0.292 0.563 ± 0.287 0.800 ± 0.122 0.547 ± 0.283 0.970 ± 0.029
Table 4: Computational time comparison (in seconds) of GAIS variants across dataset sizes. Methods include distance-based mini-batching (DM) and locality-sensitive hashing variants (SL/ML/MVML-LSH with cross-attention C). Time components: graph construction (GCT), GAT training (GTT), instance selection (IST), and machine learning model training on selected instances (MLT). Values show mean $\pm$ standard deviation. Bold indicates fastest performance per dataset category.
@@ 5.1 Comparing GAIS Variants
In medium-sized datasets, the performance gap between the original and GAIS narrows considerably. The original achieves an accuracy of 0.945, while MVML-LSH maintains robust performance with an accuracy of 0.895 and effectiveness of 0.868. The DM method demonstrates stable performance (E = 0.872) in medium datasets, outperforming several LSH variants, which suggests that its adaptive similarity threshold mechanism scales well with increasing data size. The multi-view and cross-attention variants show marginal differences in effectiveness (ranging from 0.864 to 0.869), suggesting that the additional complexity of cross-attention mechanisms might not provide substantial benefits. The reduction rates remain consistently high across all proposed methods (96-97%).
For large datasets, while MVML-LSH achieves the highest effectiveness (E = 0.819) and accuracy (AC = 0.840), its cross-attention variant CMVML-LSH demonstrates superior performance across other key metrics, including precision (PR = 0.715), recall (RE = 0.528) and F1-score (F1 = 0.563). This suggests that the cross-attention mechanism enhances the model’s ability to identify more informative instances, leading to better balanced performance. The DM approach shows competitive effectiveness (E = 0.795) while maintaining a high reduction rate (R = 0.98), demonstrating its viability as a computationally efficient alternative. The random sampling baselines consistently underperform compared to the proposed methods across all metrics, validating the importance of structured selection approaches.
@@
Table 4 presents a summary of computational time metrics for GAIS methods across different dataset sizes, comparing graph construction time (GCT), GAT (or GNN) training time (GTT), selection time (IST), and ML training time (MLT). For small datasets, all proposed methods demonstrate efficient computational performance, with GCT remaining under 2 seconds on average. The DM method shows particularly efficient graph construction (GCT = 0.05s), making it an attractive option for time-sensitive applications with small datasets. While the multi-level variants (ML-LSH, CML-LSH) show slightly higher GCT ($\approx$ 1.9s), they maintain competitive GTT around 4-5 seconds. The IST is negligible across all methods for small datasets.
@@ 5.2 Comparing Graph Timings
Figure 7(a) compares graph construction times between GAIS variants and KNN-based approaches across medium and large datasets. The DM method consistently demonstrates higher efficiency across all dataset sizes, maintaining relatively stable construction times even as dataset size increases. Among the LSH variants, SL-LSH shows the best computational efficiency, followed by ML-LSH, while MVML-LSH exhibits higher construction times due to its more complex multi-view architecture. When compared to KNN-based approaches, all GAIS variants show better scalability, particularly for larger datasets. This is most evident in datasets like SUSY, where exact KNN and FAISS exhibit exponential growth in construction time, while GAIS methods maintain more controlled growth. While approximate KNN methods like Annoy and HNSW show better scaling than exact KNN, they still require significantly more time than the GAIS variants, especially for datasets exceeding 100K instances. These results demonstrate that the proposed GAIS methods, particularly the DM approach, offer a more scalable alternative to traditional KNN-based graph construction methods.
@@ 5.3 Component Analysis
Figure 8 illustrates the relationship between mini-batch configuration and test effectiveness in the DM method for large datasets. We test configurations up to $w_{\max}$ instances per mini-batch to identify natural performance plateaus. The results demonstrate that both the number of mini-batches and instances per batch have a positive impact on effectiveness, though this improvement shows diminishing returns. As the number of mini-batches increases from 1 to 9, the effectiveness consistently improves across all instance sizes, with the most substantial gains observed in the transition from 1 to 5 mini-batches. Similarly, larger batch sizes generally yield better effectiveness, though the improvement plateaus around $7000$ instances, particularly with higher numbers of mini-batches. This plateau effect is most noticeable when using 7-9 mini-batches, suggesting that beyond these configurations, additional computational resources might not yield proportional performance benefits. These findings provide practical guidance for implementing the DM method, indicating that moderate configurations (5-7 mini-batches with $5000-7000$ instances per batch) might offer an optimal balance between computational efficiency and effectiveness. The empirical analysis demonstrates that effectiveness plateaus around $7000$ instances per mini-batch, providing direct guidance for $w_{\max}$ selection. This plateau effect validates that computational bounds can be set without sacrificing performance.
Figure 8: Impact of mini-batch configuration on test effectiveness in the distance-based mini-batch (DM) method for large datasets ($>$100K instances). Results show diminishing returns beyond 7,000 instances per mini-batch and 5-7 mini-batches, providing empirical guidance for setting the computational constraint $w_{\max}$ while maintaining performance.
Building on these mini-batch insights, we empirically validate our $w_{\text{max}}$ constraint design and address computational scalability concerns by testing performance plateaus across different sampling percentages. This analysis aims to determine the natural point where additional data sampling yields diminishing returns to justify our fixed computational bounds. Figure 9 demonstrates clear plateau patterns across dataset sizes, with performance stabilizing at low sampling percentages for larger datasets. Large datasets exhibit performance plateaus at just 5% sampling, achieving peak effectiveness with a median of 3,601 instances per mini-batch. Importantly, the correlation between dataset size and optimal mini-batch requirements for large datasets is negligible ($r=0.084$, $p=0.875$). This demonstrates that our approach scales independently of dataset size. Medium datasets plateau at 10% sampling with a median of 1,553 instances per mini-batch, while small datasets require 40% sampling with a median of 160 instances per mini-batch. These findings support that $w_{\text{max}}$ constraint captures the natural performance ceiling for large datasets while ensuring computational feasibility. The independence of optimal mini-batch size from dataset scale ($r=0.084$ for large datasets) empirically supports our $O(K\times w_{\max}^{2})$ complexity analysis.
Figure 9: Effectiveness versus sampling percentage by dataset size categories: small ($<$10K instances), medium (10K–100K instances), and large ($>$100K instances). Performance plateaus occur at 5%, 10%, and 40% sampling for large, medium, and small datasets respectively, with the ”Diminishing Returns Zone” indicating where additional sampling provides minimal benefit.
Table 5 evaluates hash bucket quality across different dataset sizes, comparing the SL- and ML-LSH approaches. The results demonstrate that ML-LSH consistently outperforms SL-LSH across most quality metrics, with particularly notable improvements in purity ($0.81$ vs $0.78$ for small datasets, $0.84$ vs $0.79$ for medium datasets) and $R@5$ performance. Most significantly, ML-LSH exhibits superior scalability characteristics: while both methods experience degraded $R@5$ as dataset size increases, ML-LSH maintains substantially higher recall rates ($0.34$ vs $0.19$ for large datasets), indicating better preservation of local neighborhood structures in large-scale settings. The multi-level approach also produces more computationally efficient bucket distributions, with consistently smaller average bucket sizes ($\bar{B}$) across all dataset categories. The strong correlation values ($\rho$) achieved by ML-LSH, particularly for large datasets ($0.64$ vs $0.59$), confirm that the hierarchical structure effectively captures meaningful similarity relationships while maintaining computational efficiency.
Table 5: Comparison of SL- and ML-LSH across dataset sizes. Results averaged across all datasets within each size category. Higher values indicate better performance for all metrics except $\bar{B}$, where smaller values indicate better computational efficiency.
Method Sep $\uparrow$ Pur $\uparrow$ $R@5$uparrow $\rho$uparrow $\bar{B}$downarrow
Small Datasets ($<10,000$)
SL 1.23 ± 0.23 0.78 ± 0.13 0.62 ± 0.29 0.45 ± 0.20 26.7 ± 24.0
ML 1.25 ± 0.25 0.81 ± 0.13 0.66 ± 0.28 0.52 ± 0.20 21.3 ± 22.6
Medium Datasets (10,000100,$0-100,00$ 100 , 000)
SL 1.21 ± 0.16 0.79 ± 0.12 0.42 ± 0.33 0.53 ± 0.22 37.2 ± 18.5
ML 1.21 ± 0.17 0.84 ± 0.10 0.54 ± 0.36 0.57 ± 0.20 28.8 ± 17.6
Large Datasets ($>100,000$)
SL 1.32 ± 0.14 0.75 ± 0.17 0.19 ± 0.20 0.59 ± 0.21 53.7 ± 28.9
ML 1.32 ± 0.15 0.79 ± 0.15 0.34 ± 0.23 0.64 ± 0.20 44.5 ± 30.0
  • Sep = Separation ratio (intra-/inter-bucket cosine similarity); Pur = Purity (class label homogeneity); $R@5$ = Recall@5 (true 5-nearest neighbors in same bucket); $\rho$ = Pearson correlation with cosine similarity; $\bar{B}$ = Average bucket size.
Figure 10 demonstrates the interpretability of our hierarchical LSH approach across two representative datasets with different manifold structures: the banana dataset with crescent-shaped clusters, and the abalone dataset with complex curved manifolds. In both cases, Level 1 identifies major structural boundaries, Level 2 refines these into coherent sub-regions, and Level 3 captures fine-grained local neighborhoods. The consistent ability to discover meaningful structures across different data geometries supports that our multi-level approach effectively approximates semantic similarity relationships while providing interpretable bucket assignments at multiple resolutions.
(a) Banana dataset
(b) Abalone dataset
Figure 10: Interpretability of hierarchical Locality-Sensitive Hashing (LSH) bucket assignments via dimensionality reduction. Our Multi-Level LSH (ML-LSH) approach partitions instances into hash buckets at multiple resolution levels to approximate similarity relationships. Each color represents a different hash bucket assignment at each level: Level 1 identifies major structural boundaries (coarse-grained cluster separation), Level 2 refines into coherent sub-regions, and Level 3 captures fine-grained local neighborhoods.
Figure 11 reveals distinct scaling patterns for hierarchical complexity in LSH-based graph construction. For multi-level analysis (Figure 11(a)), small datasets benefit most from increased hierarchy, with Level 3 achieving peak effectiveness improvement (2.69%), suggesting that fine-grained similarity capture is important when data is limited. This benefit is even more pronounced for lower-performing datasets (3.29% improvement), indicating that hierarchical structures particularly excel where simpler approaches struggle. Medium datasets show moderate but consistent gains across levels (0.85-1.6%), with Level 5 providing the highest improvement. Large datasets show smaller relative improvements (0.15-0.64%) due to already strong baseline performance from sufficient data volume, though the computational overhead of deeper hierarchies may limit additional benefits. For multi-view analysis (Figure 11(b)), the pattern is more consistent across dataset sizes: small datasets achieve the largest gains from multiple views (2.88% peak improvement with 4 views, rising to 4.08% for lower-performing datasets), as diverse similarity perspectives help overcome limited data representation. Medium and large datasets show progressively smaller but steady improvements, with lower-performing datasets consistently demonstrating larger gains than the overall average. The consistent performance plateau around 3-4 views across all dataset sizes indicates an optimal balance between representational diversity and computational efficiency, beyond which additional views provide marginal returns.
(a) Levels: Hierarchical LSH effectiveness gains
(b) Views: Multi-view LSH effectiveness gains
Figure 11: Multi-level and multi-view analysis showing relative improvement (in %) in test effectiveness for LSH-based graph construction variants. (a) Multi-level analysis: Improvement from using hierarchical levels compared to single-level baseline, across small ($<$10K instances), medium (10K–100K instances), and large ($>$100K instances) datasets. (b) Multi-view analysis: Improvement from using multiple views compared to single-view baseline. Solid lines show average improvements across all datasets in each size category, while dashed lines represent improvements for datasets performing below median effectiveness within each category to highlight challenging datasets.